A particle of mass 11.9 g and charge 89.4 muC move
Answer #1


mass of the particle, m=11.9g


charge q=89.4 uC

acceleration, a=-9.8 j m/sec^2

velocity of the partivle, V=23.6 i km/sec

===>


use,

F=q*(VXB)


m*a =89.4*10^-6*(vx i+ Vy j+ Vz k)X((Bx)i+(By)j+(Bz)k)

11.9*10^-3*(-9.8)j=89.4*10^-6*(23.6*10^3i+0j+0k)X((Bx)i+(By)j+(Bz)k)


11.9*10^-3*(-9.8)j=89.4*10^-6*((23.6*10^3By)k-(23.6*Bz)j)


11.9*10^-3*(-9.8)j=(89.4*10^-6)*(23.6*10^3By)k- (89.4*10^-6)*(23.6*10^3Bz)j

===>

comparing both sides,


Bx=0,


-(89.4*10^-6)*(23.6*10^3*Bz)= 11.9*10^-3*(-9.8)


(89.4*10^-6)*(23.6*10^3*Bz)= 11.9*9.8*10^-3


==> Bz=0.055 T


By=0

====>


magnetic field, B=(Bx)i+(By)j+(Bz)k


B=0i+0j+(0.055)k T